# RED-CT: A Systems Design Methodology for Using LLM-labeled Data to Train and Deploy Edge Classifiers for Computational Social Science

### v1

@@ Abstract
Large language models (LLMs) have enhanced our ability to rapidly analyze and classify unstructured natural language data. However, concerns regarding cost, network limitations, and security constraints have posed challenges for their integration into work processes. In this study, we adopt a systems design approach to employing LLMs as imperfect data annotators for downstream supervised learning tasks, introducing novel system intervention measures aimed at improving classification performance. Our methodology outperforms LLM-generated labels in seven of eight tests, demonstrating an effective strategy for incorporating LLMs into the design and deployment of specialized, supervised learning models present in many industry use cases.
RED-CT: A Systems Design Methodology for Using LLM-labeled Data to Train and Deploy Edge Classifiers for Computational Social Science
@@ 1 Introduction
Figure 1: RED-CT design which allows LLM-like capabilities for CSS tasks deployed in edge environments.
@@
To balance the flexibility associated with LLMs and advantages of supervised models, we propose Rapid Edge Deployment for CSS Tasks (RED-CT). RED-CT is a system that pairs LLM generated data labels with minimal human-annotated labels to deploy classifiers to edge environments fast. We define the edge environment as time- and / or resource-limited situations where users need to interface with NLP solutions. Additionally, the edge environment may be disconnected from the internet for security or privacy purposes or in crisis response settings where connection to the internet is either impracticable or unreliable.
In this paper, we introduce RED-CT and propose a novel confidence-informed sampling method to select LLM-labeled data for human annotation. In addition, we present a simple method to generate soft labels from LLM predictions to use during edge classifier training. We evaluate RED-CT with confidence-informed sampling and learning on soft labels across four CSS tasks: stance detection, misinformation identification, humor detection, and ideology detection. We further evaluate the proposed approach with two different common data-labeling prompting schemes and across three different sizes of distilled models. Our results show that it is possible to approximate or outperform LLMs on CSS tasks with minimal human data labeling ( $\leavevmode\nobreak\ 10\%$ of dataset) in the distillation of edge models.
@@ 2 Related Works
Recent work has demonstrated that LLMs can perform well across various classification tasks within CSS Cruickshank and Ng (2024); Zhu et al. (2023). Ziems et al. (2024) provides best practices for prompting and benchmarks performance for a variety of CSS tasks across several LLMs. LLM-based classification methods work better with out-of-domain data due to the LLMs strong zero-shot classification capacity. However, these methods also require substantial resources and cannot scale, in terms of cost or compute time, to large CSS datasets. For example, just labeling the SemEval2016 dataset Mohammad et al. (2016) (2,814 data points) with GPT-4 could cost over \$30 USD. Additionally, ongoing research has found that LLMs still usually perform worse than in-domain supervised models at CSS tasks Cruickshank and Ng (2024); Ziems et al. (2024).
@@
Researchers study LLM distillation Xu et al. (2024) techniques to reduce model size and cost. These methods vary considerably in their use of LLMs. Some studies have focused on generating artificial data with LLMs useful for distilling small classification models Ye et al. (2022b, a); Gao et al. (2022); Meng et al. (2023). Other works have explored few-shot prompting and active learning mechanisms, combined with LLMs for data labeling Wang et al. (2021); Zhang et al. (2023); Hsieh et al. (2023). Many of these methods often require human intervention to filter low-quality data or LLM-generated rationales for labels which can be unreliable Huang et al. (2023). Other works focus on reducing bias Egami et al. (2023) without focusing on downstream classification performance Wang et al. (2021). Pangakis and Wolken (2024) assess supervised classifiers performance on LLM generated labels, but do not offer a systems approach or intervention measures to improve downstream classification. None of the prior works attempt to integrate additional uncertainty information from LLMs into human intervention and model distillation.
@@ 3.2.1 Incorporating Confidence Informed Expert Labels
Randomly selecting samples for SME labeling within a bounded time or up to a certain percentage can improve edge model performance but may introduce inefficiencies where SME’s analyze sample data in which the LLM is confident it has labeled correctly. To optimize sampling for SME analysis, we devise a confidence-based metric to identify examples where LLM labeling is less reliable. We estimate label confidence scores using the absolute difference between the maximum label log-probability and second-highest label log probability. To apply the confidence score, we stratify by each LLM-labeled class and sample the bottom $p$ percentile.
@@
Figure 2: The distribution of confidence scores for examples labeled correctly and incorrectly using gpt-3.5-turbo zero-shot stance classification.
@@ 3.2.2 Learning on Soft Labels
GPT-3.5 Turbo Random Base RS 10% CI LW 10%
@@
Table 1: Zero-Shot LLM performance (weighted f1 score) compared to edge model performance. Random are dummy models predicting on a uniformed distribution, base edge models are trained without system interventions, RS 10% edge models are trained with 10% randomly sampled expert examples, and CI LW 10% is 10% confidence-informed sampling and learning with label weights.
@@ 4 System Implementation and Experiment Design
For edge classifiers, we test three flavors of BERT: ’Distil-BERT’, ’RoBERTa’, and ‘RoBERTa-Large’ Devlin et al. (2019); Liu et al. (2019). These models vary in size, allowing us to assess performance across model compute requirements. For each BERT model, we evaluate the effects of system intervention measures. The system settings we tested included a base classifier trained with no system interventions directly on the LLM labels, a classifier trained on soft labels (LW), a classifier trained on 10 percent randomly selected expert labeled data (RS 10%), a classifier trained on confidence-informed sampling (CI 10%), and a classifier trained with all system intervention measures (CI LW 10%). We train five classifiers on each LLM-labeled dataset and report the averages across each. For each edge model, we do full fine-tuning (i.e., unfreeze all model weights) from pre-trained models, but note that this process can be done with any type of fine-tuning or training a model with initialized weights.
@@ 5 Results
Table 1 presents the high-level results across our four chosen CSS tasks using RoBERTa-L. Figure 3 is a more detailed analysis of the implementation of our system methodology in the stance detection task, including varying the type of BERT model in all combinations of system intervention strategies. A key takeaway is that through our methodology and associated system intervention measures, we were able to outperform LLM-labeled data in 7 of the 8 tested tasks, while approximating it in an additional task. Additionally, in GPT labeled data, we had an average improvement of 6.5% over the base classifier. In Annex B, we have included additional results analysis, including varying the percentage of expert labels in Figure 4 and a full table for each stance detection result in Table 2.
@@ 6 Conclusions
Using a systems methodology, we have successfully replicated LLM performance in an edge environment on CSS tasks in a zero-shot setting. Our system does require expert-in-the-loop data labeling for a small portion of the data (10% or less to obtain highly performant small models), but allows for model deployment into environments where LLM access is restricted for either cost, security, or latency reasons. Our results hold across a number of labeling prompts and distilled models, showing the generalizability of our approach. Through our system design, we believe that we have found an improved mechanism for a few-shot sample selection in a few-shot labeling on LLMs.
6.1 Future Work
A question left unanswered by this work is how different prompting techniques influence logit values and resulting confidence scores. Integration of confidence scores for few-shot selection in downstream classification tasks could have broader implications in machine learning, especially in more decentralized approaches such as federated learning. Another question that follows is are there better logit transformation techniques for resulting confidence-informed sampling. While effective, our transformation technique was relatively simple and we did not test additional transformation mechanisms. Lastly, we only tested our systems methodology on a subset of CSS tasks, but our intervention measures can likely be applied to several other domains in downstream classification where LLMs are used as imperfect annotators such as image classification or spam detection.
6.2 Limitations
A limitation of the proposed methodology is that it is only as good as the labeled data and the LLM. In higher performing LLMs, our system intervention mechanisms worked better as the confidence and weight scores were more meaningful. Our system also relies on windows of connectivity for both data transfer and model deployment. The proposed method is designed for specific tasks and deployed models only perform well on data on which they were trained; a model trained on stances about election rumours may not transfer well to stances on sports teams, for example. Lastly, our system is not completely autonomous and does benefit greatly from even minimal human participation.
Ethics and Availability
To maintain transparency and reproducibility, we used public datasets and models. The code is available at https://anonymous.4open.science/r/REDCT-C6A5/. The SemEval2016 dataset is available at https://www.saifmohammad.com/WebPages/StanceDataset.htm, the Misinfo Reaction Frames dataset is available at https://github.com/skgabriel/mrf-modeling, and the Reddit Jokes Database is available at https://github.com/orionw/RedditHumorDetection/tree/master/data. The IBC Dataset is available at https://github.com/SALT-NLP/LLMs_for_CSS/tree/main/css_data/ibc. Each dataset contains a collection of social media statements.
We query GPT-3.5 through OpenAIs public API (https://platform.openai.com/docs/api-reference). The 7B parameter Mistral model is available at https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2. The three flavors of BERT used distilBERT, RoBERTa, and RoBERTa-L are available at https://huggingface.co/docs/transformers/en/model_doc/distilbert, https://huggingface.co/FacebookAI/roberta-base, and https://huggingface.co/FacebookAI/roberta-large respectively. We used each model in accordance with their respective terms of use.
@@ Appendix B Stance Classification Additional Results
Figure 4: Varying the number of expert labels included amongst the LLM labels in the training process for DistilBERT and RoBERTa-L. RS implies randomly sampled expert labels for the training process and CI LW implies confidence informed sampling with label weighted training. Blue corresponds to the Mistral-7B-Instruct-2.0 LLM labeler and green corresponds to the GPT-3.5 LLM labeler. The horizontal dashed lines represent the zero-shot accuracy of each LLM.
@@
Table 2: F1 scores on SemEval2016. Edge classifier variants: ’Base’ trained on LLM labels directly, LW trained with label weighting, RS 10% trained with 10% randomly sampled expert labels, CI 10% trained with 10% confidence informed expert labels, and CI LW 10% trained with 10% confidence informed expert labels and labeling weighting.

### v2

# RED-CT: A Systems Design Methodology for Using LLM-labeled Data to Train and Deploy Edge Linguistic Classifiers

@@ Abstract
Large language models (LLMs) have enhanced our ability to rapidly analyze and classify unstructured natural language data. However, concerns regarding cost, network limitations, and security constraints have posed challenges for their integration into industry processes. In this study, we adopt a systems design approach to employing LLMs as imperfect data annotators for downstream supervised learning tasks, introducing system intervention measures aimed at improving classification performance. Our methodology outperforms LLM-generated labels in six of eight tests and base classifiers in all tests, demonstrating an effective strategy for incorporating LLMs into the design and deployment of specialized, supervised learning models present in many industry use cases.
RED-CT: A Systems Design Methodology for Using LLM-labeled Data to Train and Deploy Edge Linguistic Classifiers
David Farr 1,2, Nico Manzonelli 2, Iain Cruickshank 3, Jevin West1 1University of Washington, 2Army Cyber Technology and Innovation Center, 3Carnegie Mellon University Correspondence: dtfarr@uw.edu
@@ 1 Introduction
Figure 1: RED-CT design which allows LLM-like capabilities for NLP tasks deployed in edge environments.
@@
To balance the flexibility associated with LLMs and advantages of supervised models, we propose Rapid Edge Deployment for CSS Tasks (RED-CT). RED-CT is a system that integrates traditional techniques from active learning such as confidence measurements and soft labels to pair LLM generated data labels with minimal selected human-annotated labels to deploy classifiers to edge environments fast. We define the edge environment as time- and / or resource-limited situations where users need to interface with NLP solutions. Additionally, the edge environment may be disconnected from the internet for security or privacy purposes or in crisis response settings where connection to the internet is either impracticable or unreliable.
In this paper, we introduce RED-CT and propose a confidence-informed sampling method to select LLM-labeled data for human annotation. In addition, we present a simple method to generate soft labels from LLM predictions to use during edge classifier training. We evaluate RED-CT with confidence-informed sampling and learning on soft labels across four CSS tasks: stance detection, misinformation identification, humor detection, and ideology detection. We further evaluate the proposed approach with two different common data-labeling prompting schemes and across three different sizes of distilled models. Our results show that it is possible to approximate or outperform LLMs on CSS tasks with minimal human data labeling (${}10\%$ of dataset) in the distillation of edge models.
@@ 2 Related Works
Recent work has demonstrated that LLMs can perform well across various classification tasks within CSS Cruickshank and Ng (2024); Zhu et al. (2023). Ziems et al. (2024) provides best practices for prompting and benchmarks performance for a variety of CSS tasks across several LLMs. LLM-based classification methods work better with out-of-domain data due to the LLMs strong zero-shot classification capacity. However, these methods also require substantial resources and cannot scale, in terms of cost or compute time, to large CSS datasets. For example, just labeling the SemEval2016 dataset Mohammad et al. (2016) (2,814 data points) with GPT-4 could cost over \$30 USD. Additionally, ongoing research has found that LLMs still usually perform worse than in-domain supervised models at CSS tasks Cruickshank and Ng (2024); Ziems et al. (2024). Tan et al. (2024) provide a survey paper of research using large language models for data annotation, including model distillation as a task. Some related works included show using synthetic generated data from larger LLMs to train smaller LLMsWang et al. (2023) and Huang et al. (2022) which demonstrate LLMs can improve performance through self-annotation and subsequent fine-tuning based on self-annotated data. Further related work by Wang et al. (2024) deploys an external verifier model to select samples LLMs are unlikely to classify correctly and routes them to human labelers for increased performance. Such previous work differs in data sampling methods, resource requirements, and distillation methodology.
@@
Researchers study LLM distillation techniques Xu et al. (2024) to reduce model size and cost. These methods vary considerably in their use of LLMs. Some studies have focused on generating artificial data with LLMs useful for distilling small classification models Ye et al. (2022b, a); Gao et al. (2022); Meng et al. (2023). Other works have explored few-shot prompting and active learning mechanisms, combined with LLMs for data labeling Wang et al. (2021); Zhang et al. (2023); Hsieh et al. (2023). Many of these methods often require human intervention to filter low-quality data or LLM-generated rationales for labels which can be unreliable Huang et al. (2023). Other works focus on reducing bias Egami et al. (2023) without focusing on downstream classification performance Wang et al. (2021). Pangakis and Wolken (2024) assess supervised classifiers performance on LLM generated labels, but do not offer a systems approach or intervention measures to improve downstream classification. None of the prior works attempt to integrate additional uncertainty information from LLMs into human intervention and model distillation.
@@ 3.2.1 Incorporating Confidence Informed Expert Labels
Randomly selecting samples for SME labeling within a bounded time or up to a certain percentage can improve edge model performance but may introduce inefficiencies where SME’s analyze sample data in which the LLM is confident it has labeled correctly. To optimize sampling for SME analysis, we devise a confidence-based metric to identify examples where LLM labeling is less reliable.
The confidence score is defined as the absolute difference between the highest token label log probability and the second-highest token label log probability within this constrained set of expected tokens. Let $\mathcal{T}$ represent the set of given tokens, and $P(t)$ denote the distribution of log probabilities across each token $t\in\mathcal{T}$. The confidence score, denoted as $C$, is then computed using the formula
[…]
where $t^{*}$ is the token corresponding to the highest probability $\max_{t\in\mathcal{T}}P(t)$. To apply the confidence score, we stratify by each LLM-labeled class and sample the bottom $p$ percentile.
@@
Figure 2: The distribution of confidence scores for examples labeled correctly and incorrectly using gpt-3.5-turbo zero-shot stance classification. The distributions are overlaid as opposed to stacked.
@@ 3.2.2 Learning on Soft Labels
GPT-3.5 Turbo Random Base RS 10% CI SL 10%
@@
Table 1: Zero-Shot LLM performance (weighted f1 score) compared to edge model performance. Random are dummy models predicting on a uniformed distribution, base edge models are trained without system interventions, RS 10% edge models are trained with 10% randomly sampled expert examples, and CI SW 10% is 10% confidence-informed sampling and learning with label weights. Results that out-performed the enterprise LLM are bolded.
@@ 4 System Implementation and Experiment Design
For edge classifiers, we test three flavors of BERT: ’Distil-BERT’, ’RoBERTa’, and ‘RoBERTa-Large’ Devlin et al. (2019); Liu et al. (2019). These models vary in size, allowing us to assess performance across model compute requirements. For each BERT model, we evaluate the effects of system intervention measures. The system settings we tested included a base classifier trained with no system interventions directly on the LLM labels, a classifier trained on soft labels (SL), a classifier trained on 10 percent randomly selected expert labeled data (RS 10%), a classifier trained on confidence-informed sampling (CI 10%), and a classifier trained with all system intervention measures (CI SL 10%). We train five classifiers on each LLM-labeled dataset and report the averages across each. For each edge model, we do full fine-tuning (i.e., unfreeze all model weights) from pre-trained models, but note that this process can be done with any type of fine-tuning or training a model with initialized weights.
@@ 5 Results
Table 1 presents the high-level results across our four chosen CSS tasks using RoBERTa-L. Figure 3 is a more detailed analysis of the implementation of our system methodology in the stance detection task, including varying the type of BERT model in all combinations of system intervention strategies. A key takeaway is that through our methodology and associated system intervention measures, we were able to outperform LLM-labeled data in 6 of the 8 tested tasks, while approximating it in an additional task. Additionally, in GPT labeled data, we had an average improvement of 6.75% over the base classifier and we out performed the base and normal sampling techniques in 100 percent of tasks. In Annex B, we have included additional results analysis, including varying the percentage of expert labels in Figure 4 and a full table for each stance detection result in Table 2.
@@ 6 Conclusions
In this work, we successfully replicated LLM performance in an edge environment on computational social science tasks using a systems methodology. Our approach, which integrates expert-in-the-loop data labeling for a small portion of the data (10% or less), enables the deployment of highly performant small models in environments where LLM access is restricted due to cost, security, or latency concerns.
Our results demonstrate generalizability across various labeling prompts and distilled models, providing a flexible and scalable solution. This methodology offers a practical mechanism to reduce labeling costs and dependence on large LLMs while improving performance and data annotation throughput, even in resource-constrained settings with minimal human intervention.
@@ Appendix B Stance Classification Additional Results
Figure 4: Varying the number of expert labels included amongst the LLM labels in the training process for DistilBERT and RoBERTa-L. RS implies randomly sampled expert labels for the training process and CI SL implies confidence informed sampling with label weighted training. Blue corresponds to the Mistral-7B-Instruct-2.0 LLM labeler and green corresponds to the GPT-3.5 LLM labeler. The horizontal dashed lines represent the zero-shot accuracy of each LLM.
@@
Table 2: F1 scores on SemEval2016. Edge classifier variants: ’Base’ trained on LLM labels directly, SL trained with label weighting, RS 10% trained with 10% randomly sampled expert labels, CI 10% trained with 10% confidence informed expert labels, and CI SL 10% trained with 10% confidence informed expert labels and labeling weighting.